3) The success rate of the seller is 0.1. He can sell (success) or not. What...

3)

The success rate of the seller is 0.1. He can sell (success) or not. What is the probability of one success in five customers.
Answer #1

solution

Given that ,

p = 0.1

1 - p = 1 - 0.1 = 0.9

n =5

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X =1 ) = (5 C 1) * 0.11 * (0.90)4

probability =0.3281